OpenStudy (firejay5):

Show work!!! :D Medal will be rewarded! :D 27. Basketball: Kiana has made 9 of 19 free throws so far this season. Her goal is to make 60% of her free throws. If Kiana makes her next x free throws in a row, the function: f(x) = x + 9 over x + 19 represents Kiana's new ration of free throws made. How many successful free throws in a row will raise Kiana's percent made to 60%.

OpenStudy (mertsj):

\[\frac{x+9}{x+19}=\frac{60}{100}\]

OpenStudy (mertsj):

Solve that.

OpenStudy (firejay5):

solve for x correct @Mertsj

OpenStudy (mertsj):

yes. I would cross multiply